OpenStudy (anonymous):

Find the range of the function. f(x) = (x – 2)^2 + 2 (2 points) All real numbers y ≥ 0 y > 2 y ≥ 2

OpenStudy (anonymous):

hint \[(x-2)^2 \ge 0\]

OpenStudy (anonymous):

can you show me how you got that

OpenStudy (anonymous):

u dont need to get that from somewhere square of every real number is greater than or equal to zero

OpenStudy (anonymous):

can u tel me what is the final answer plz?

OpenStudy (anonymous):

y ≥ 0

OpenStudy (anonymous):

\[(x-2)^2\ge 0\]add 2 to both sides\[(x-2)^2+2\ge 2\]

OpenStudy (anonymous):

y ≥ 2?